OpenStudy (anonymous):

Separable Equations \[\frac{dy}{dx}=\frac{\sqrt{x}}{e^{y}}\] \[\int e^{y}dy=\int \sqrt{x}dx \] \[\frac{e^{y}}{2x}=y\] Yes?

OpenStudy (anonymous):

Nope. \[\int e^{y}dy \neq \int \sqrt{x}dx\] \[\int e^{y}dy = e^y + C\] \[\int \sqrt{x} dx = \int x^{1/2} dx = \frac{x^{1/2+1}}{1/2+1} + C = \frac{2}{3}x^{3/2} + C\]

OpenStudy (anonymous):

Your second line of logic is false.

myininaya (myininaya):

wait? Why is the second line that he has is bad?

OpenStudy (anonymous):

Disagree?

OpenStudy (anonymous):

To solve for Y, you need to equate both answers and take the natural log of both sides.

myininaya (myininaya):

Well we want to separate the variables \[\frac{dy}{dx}=\frac{\sqrt{x}}{e^y}\] Multiply e^y on both sides Multiply dx on both sides \[e^y dy =\sqrt{x} dx\] Now integrate both sides

myininaya (myininaya):

\[\int\limits_{}^{}e^y dy= \int\limits_{}^{}\sqrt{x} dx\]

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

is my integration correction though?

myininaya (myininaya):

\[\int\limits_{}^{}e^y dy = \int\limits_{}^{}x^\frac{1}{2} dx\]

myininaya (myininaya):

Hey @MathSofiya agent showed you how to integrate the right side

myininaya (myininaya):

and the left side

OpenStudy (anonymous):

ok...so I did it wrong, lolz

OpenStudy (anonymous):

After integrating, your answer should look like: e^y = 2/3x^3/2 + c. Then, take the natural log of both sides to get Y.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

\[\ln|e^y| = ln|(2/3)x^{3/2} + C|\] \[y = ln|(2/3)x^{3/2}| + C\] Although why we don't have to rename C is a technical mystery I've never quite understood. I stand corrected @myininaya , it's the 3\(^{rd}\) line, not the 2\(^{nd}\). :-)

OpenStudy (anonymous):

thanks agentx5

OpenStudy (anonymous):

why the absolute value?

myininaya (myininaya):

@agentx5 lol okay

OpenStudy (anonymous):

Because you can't rule out if the result of what's inside is negative for some reason, the domain of logarithms between this: 0<x<\(\infty\) I'll leave you to think about why ;-) PS: If there's some sort of conditions we can use to know (2/3)x^(3/2) is always positive then we can drop the absolute value bars

OpenStudy (anonymous):

fair enough

OpenStudy (anonymous):

:)